# ComFace: Facial Representation Learning with Synthetic Data for Comparing Faces

### v1

@@ Abstract
Keywords:
Facial representation learning Intra-personal facial changes Learning from synthetic data Transfer learning Contrastive learning
@@ 2 Related Work
Facial Representation Learning: Transfer learning in face analysis tasks commonly depends on pre-training using ImageNet [15, 2, 45, 37] and large face recognition datasets [57, 58, 48, 32, 52]. Recently, facial representation learning (FRL) has been studied as a pre-training method for face analysis tasks [6, 72, 39]. Bulat et al. [6] investigated pre-training strategies and datasets for several face analysis tasks. They proposed a SwAV [8]-based self-supervised pre-training method trained on unlabeled face images. Zheng et al. [72] proposed a weakly-supervised method called FaRL using face image-text pairs. FaRL adopts image-text contrastive learning to obtain high-level semantic meaning and masked image modeling to capture low-level information. Liu et al. [39] proposed a pose-disentangled contrastive learning (PCL) method for general self-supervised facial representation. PCL disentangles the learning on pose-related features and pose-unrelated facial features. Different from the FRL methods described above, our ComFace focuses on FRL regarding intra-personal facial changes.
@@ 3.1 Synthetic Face Images
Our method utilizes synthetic face images generated by StyleGAN [29, 27, 30, 28] for representation learning. StyleGAN is one of the most popular generative models for face synthesis, demonstrating impressive performance in image generation, inversion, and manipulation [4]. StyleGAN generators synthesize exceedingly realistic images and enable editing [56, 11, 1, 3, 38, 49]. InterFaceGAN [56] is a major framework for face editing. It semantically edits faces by interpreting latent semantics learned by StyleGAN. Specifically, InterFaceGAN can easily edit latent code $\bm{w}$ in high dimensional space $\mathcal{W}$ to manipulate the attributes (e.g., weight, age, and gender) of a synthesized image as $\bm{w}_{edit}=\bm{w}+\alpha\cdot\bm{n}_{m}$, where $\bm{w}_{edit}$ is the edited latent code, $\bm{n}_{m}$ is a normal vector to manipulate for attribute $m$, and $\alpha$ is a parameter that controls the intensity of the manipulation. It will make the synthesis look more positive for each attribute when $\alpha>0$ (e.g., get fat when the attribute $m$ is weight), and $\alpha<0$ will make the synthesis look more negative (e.g., slim down when the attribute $m$ is weight). As attributes to manipulate synthetic faces, we use weight [50], age [56], smile [56], and the 40 attributes included in the CelebA dataset [40]. Examples of synthetic images manipulated when the attribute $m$ is weight are shown on the left side of Fig. 1. We synthesize a new person’s face by randomly calculating the latent code $\bm{w}$, and generate images of intra-personal face changes by setting multiple $\alpha$.
@@ 3.2 FRL with Synthetic Face Images
[…]
@@
Intra-personal Learning: We acquire feature representations of intra-personal facial changes. The synthetic image $\bm{y}_{i}$ is converted into the image $\tilde{\bm{y}_{i}}$ by data augmentation. Then, $\tilde{\bm{y}_{i}}$ is input to $f(\cdot)$, which outputs the vector $\bm{h}_{\tilde{\bm{y}}_{i}}=f(\tilde{\bm{y}}_{i})$. We next calculate the difference vector $\bm{h}_{\tilde{\bm{y}}_{i}}-\bm{h}_{\tilde{\bm{x}}_{i}}$ between $\bm{h}_{\tilde{\bm{x}}_{i}}$ and $\bm{h}_{\tilde{\bm{y}}_{i}}$ to capture intra-personal facial changes. Furthermore, the difference vector is input to ReLU and a linear layer $l(\cdot)$, which outputs the distance of intra-personal facial changes $d_{i}$. ComFace learns how much the intra-personal face has changed on the basis of the mean squared error (MSE) loss: $\mathcal{L}_{intra}=\frac{1}{N}\sum_{i=1}^{N}(d_{i}-|\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}|)^{2}$, where $|\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}|$ is the ground truth (GT) for the distance of intra-personal facial changes. Since the direction of facial change depends on the attribute, we use $|\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}|$ instead of $\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}$ as the intra-personal GT. The network structure in intra-personal learning is similar to the Siamese network in Ref. [55].
@@ 3.3 Curriculum Learning of Intra-personal Facial Changes
In intra-personal learning, the difficulty level of learning tasks depends on $\alpha$. Specifically, when $\alpha=0.0$ and $5.0$, there are significant differences in the facial changes, so it is easy to identify them (see the left side of Fig. 1). On the other hand, when $\alpha=0.0$ and $2.0$, the facial changes are small, so it is difficult to identify them (see the left side of Fig. 1). ComFace provides effective representation learning by increasing the difficulty level of learning tasks on the basis of curriculum learning. Curriculum learning [5], which mimics the human learning behavior of starting with simple tasks and gradually learning more complex concepts, has also been shown to be effective in video representation learning [54]. ComFace gradually decreases the distance between intra-personal facial changes $|\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}|$ during training to acquire more effective feature representations. Specifically, suppose that $S$ is the range of the distance of facial changes $|\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}|$ when sampling the pairs of $\bm{x}_{i}$ and $\bm{y}_{i}$ used for training, and we gradually decrease $S$ according to the number of epochs as follows:
@@ 4.1 Setup for FRL
Synthetic Face Images: We utilize synthetic face images generated by StyleGAN [29] and StyleGAN3 [28]. The reason for using two different StyleGANs is to employ the attributes for face manipulation provided in each StyleGAN. In StyleGAN, we use the attributes that vary weight provided by Ref. [50] and the attributes that vary age and smile provided by InterFaceGAN [56]. In StyleGAN3, we employ the 40 attributes included in the CelebA dataset [40] provided by Ref. [4]. The 40 attributes contain various face-related factors such as Big Nose, Bags Under Eyes, and Pale Skin. With these 43 attributes, we use synthetic face images that change according to a wide range of face attributes. We generate synthetic face images with 250,000 identities using StyleGANs. For each identity, synthetic images for each attribute with $\alpha=\{-5.0,-4.9,\cdots,0.0,\cdots,4.9,5.0\}$ are generated (see Section 3.1). As a result, the total number of synthetic face images used for FRL is 35 million (M). For details on the composition of synthetic images, please refer to the supplementary material.
Details for FRL: For FRL with synthetic images, we use 90% of all identities as training data and 10% as validation data. As a backbone $f(\cdot)$, we use ResNet50 [23], which is commonly used for general pre-training [13, 7] and representation learning [9, 22, 6]. Since ResNet50 is better than ViT [17] for ComFace, we employ ResNet50 (see Section 4.6 for details). Synthetic images are resized to 224$\times$224 and the temperature parameter $\tau$ is set to 0.1. Our model is trained from scratch with randomly initialized weights. We run the training for 12 epochs with batch size 1024 on 32 NVIDIA A100 GPUs ($\sim$22 hours training). As parameters for curriculum learning, we set $S_{max}=10,\;t\in\{2,3,4\},\;e_{1}=3,\;e_{2}=6,\;e_{3}=9,\;e_{4}=12$. For each epoch, we randomly sample $\bm{x}_{i}$ and $\bm{y}_{i}$ pairs according to the range $S$ in curriculum learning and construct mini-batches. Adam [31] optimizer is used, and the learning rate is initialized as 4e-4 and halved in 10 epochs. See the supplemental material for other details.
@@ 4.2 Setup for Downstream Tasks
Facial Expression Change Dataset: We use the public dataset DISFA [42, 41], which contains facial videos of 27 subjects (12 women and 15 men) while they watch a 4-minute video intended to elicit a range of facial expressions. For each subject, 4845 video frames were recorded, and the action unit (AU) intensity was annotated for each frame with six levels from 0 (not present) to 5 (maximum intensity) for several AUs. We use AU6 (Cheek Raiser) and AU12 (Lip Corner Puller), which contain frames with a high AU intensity, and we extract frames with an intensity from 1 to 5. We estimate the intensity changes of AU6 and AU12 as facial expression changes.
@@
Details for Downstream Tasks: For each dataset, we perform a four-fold cross-validation that splits subjects between the training and test data. Hence, we evaluate the generalization performance for new subjects not included in the training data. We use 10% of the subjects in the training data as validation data. For both training and testing, the estimation of intra-personal facial changes is performed using two randomly sampled face images within an individual (i.e., approximately the same number of samples in the direction of increasing or decreasing changes). Note that the sampled pairs are identical across methods. For transfer learning, we use two types of evaluation, linear evaluation (i.e., backbone $f(\cdot)$ is frozen and linear layer $l(\cdot)$ is trained from scratch) and fine-tuning (i.e., all layers in $f(\cdot)$ and $l(\cdot)$ are trained from pre-trained weights), as in the previous representation learning studies [9, 22, 6]. See the supplemental material for other details.
@@ 4.3 Comparative Methods and Evaluation Metrics
We are curious about the following questions: Does ComFace successfully acquire facial representations that capture intra-personal facial changes? How is the transfer performance of representation learning using synthetic images against real images? To answer these questions, we compare ComFace with four types of methods: scratch (random initialization), general pre-training, visual representation learning, and facial representation learning. We use a supervised method using ImageNet [13] and a face recognition method using VGGFace2 [7] for general pre-training and four popular methods for visual representation learning. In the above methods, we use ResNet50, which is the same backbone as ours for fair comparison. For facial representation learning, we use Bulat et al. [6], FaRL [72], and PCL [39]. Table 1 summarizes the training datasets, training scales, training sources, and backbones for all comparative and proposed methods.
@@
Table 1: Training datasets, training scales, training sources, and backbones for all comparative and proposed methods.
Method Dataset Training Scale Training Source Backbone
Scratch - - - ResNet50
General Pre-training:
ImageNet [13] ImageNet 1.28M Images+Human labels ResNet50
VGGFace2 [7] VGGFace2 3.31M Images+Human labels ResNet50
Visual Representation Learning:
SimCLR [9] ImageNet 1.28M Images ResNet50
MoCo v2 [22] ImageNet 1.28M Images ResNet50
SwAV [8] ImageNet 1.28M Images ResNet50
Barlow Twins [67] ImageNet 1.28M Images ResNet50
Facial Representation Learning:
Bulat et al. [6] VGGFace 3.4M Face images ResNet50
FaRL [72] LAION-FACE [72] 20M Face images+Text ViT-B/16
PCL [39] VoxCeleb1 [44]+VoxCeleb2 [10] unknown Face images 16-layer CNN
ComFace (Ours) Synthetic data 35M Synthetic face images+Intensity $\alpha$ ResNet50
@@ 4.4 Main Results
We also compare ComFace with a recent visual representation learning method using synthetic data, StableRep [60]. Table 4 shows the results of weight change estimation in representation learning methods using synthetic data. Although our model has a smaller training scale than StableRep, ComFace still has superior transfer performance. We expect that this is because the representations learned using synthetic face images from StyleGANs are more suitable for the estimation of intra-personal facial changes than those learned using general images from Stable Diffusion [53].
Table 4: Results of weight change estimation in representation learning methods using synthetic data. Results are evaluated in fine-tuning. Training scales, generative models for synthesis, and backbones are described.
[…]
StableRep [60] 100M Stable Diffusion ViT-B/16 1.434 0.725 86.5 1.439 0.799 93.9 1.715 0.809 93.3
ComFace (Ours) 35M StyleGANs ResNet50 1.394 0.750 88.6 1.523 0.801 96.3 1.668 0.819 93.8
@@
Table 6 shows the results of estimating age change in fine-tuning. The best models from the four types of comparative methods are listed (see the supplemental material for full versions). The table indicates that ComFace has a higher transfer performance for age change than all other methods. We find that our method can be successfully adapted to the task related to long time periods (1 $\sim$) of temporal changes in the face.
@@ 4.5 Ablation Study
Table 8: Performance for several backbones. Results are evaluated in fine-tuning.
Backbone AU6 AU12 Edema-A Edema-B Age
ResNet18 0.658 0.826 87.2 93.2 0.870
ResNet50 0.663 0.831 88.6 96.3 0.870
ResNet101 0.660 0.820 86.4 93.6 0.853
ViT-B/16 0.609 0.825 81.8 93.9 0.843
4.6 Backbones
Table 8 compares performance in several backbones for ComFace. ResNet50 is more suitable for ComFace than ViT and other scales of ResNets. This could be because the dataset size for downstream tasks is not large and a medium-scale network performs better.

### v3

@@ 2 Related Work
Facial Representation Learning: Transfer learning in face analysis tasks commonly depends on pre-training using ImageNet [15, 2, 45, 38] and large face recognition datasets [53, 54, 47, 33, 51]. Recently, facial representation learning (FRL) has been studied as a pre-training method for face analysis tasks [6, 70, 40]. Bulat et al. [6] investigated pre-training strategies and datasets for several face analysis tasks. Zheng et al. [70] proposed a weakly-supervised method called FaRL using face image-text pairs. Liu et al. [40] proposed a pose-disentangled contrastive learning (PCL) method for general self-supervised facial representation. Different from the FRL methods described above, our ComFace focuses on FRL regarding intra-personal facial changes. Inter- and intra-personal influences on human interactions have been studied previously [31, 63]. These methods simply handle inter- and intra-personal multimodal signals ($e.g.,$ video [31], audio [31, 63], facial gestures [63]) to consider human-human interactions for specific applications such as modeling affect dynamics and adapted behavior synthesis. Our work differs from previous work [31, 63] in that our work is a comprehensive representation learning study for downstream tasks that capture intra-personal facial changes.
@@ 3.1 Synthetic Face Images
Our method utilizes synthetic face images generated by StyleGANs [29, 27, 30, 28] for representation learning. StyleGAN is one of the most popular generative models for face synthesis, demonstrating impressive performance in image generation, inversion, and manipulation [4]. StyleGAN generators synthesize exceedingly realistic images and enable editing [52, 11, 1, 3, 39, 48]. InterFaceGAN [52] is a major framework for face editing. It semantically edits faces by interpreting latent semantics learned by StyleGAN. Specifically, InterFaceGAN can easily edit latent code $\bm{w}$ in high dimensional space $\mathcal{W}$ to manipulate the attributes (e.g., weight, age, and gender) of a synthesized image as $\bm{w}_{edit}=\bm{w}+\alpha\cdot\bm{n}_{m}$, where $\bm{w}_{edit}$ is the edited latent code, $\bm{n}_{m}$ is a normal vector to manipulate for attribute $m$, and $\alpha$ is a parameter that controls the intensity of the manipulation. It will make the synthesis look more positive for each attribute when $\alpha>0$ (e.g., get fat when the attribute $m$ is weight), and $\alpha<0$ will make the synthesis look more negative (e.g., slim down when the attribute $m$ is weight). As attributes to manipulate synthetic faces, we use weight [49], age [52], smile [52], and the 40 attributes included in the CelebA dataset [41]. Examples of synthetic images manipulated when the attribute $m$ is weight are shown on the left side of Fig. 1. We synthesize a new person’s face by randomly calculating the latent code $\bm{w}$, and generate images of intra-personal face changes by setting multiple $\alpha$.
@@ 3.2 FRL with Synthetic Face Images
Inter-personal Learning: We use contrastive learning to acquire feature representations of facial differences between individuals. Contrastive learning is a kind of self-supervised learning that performs representation learning without supervised labels [9, 22], attracting positive pairs (e.g., the same images with different data augmentation) and pulling negative pairs (e.g., different images) away. ComFace performs contrastive learning based on SimCLR [9]. First, the synthetic image $\bm{x}_{i}$ is converted into two images $\tilde{\bm{x}}_{i}$ and $\tilde{\bm{x}}_{i+N}$ by data augmentation consisting of horizontal flip, color jitter, grayscale conversion, and random crop. Since $\tilde{\bm{x}}_{i}$ and $\tilde{\bm{x}}_{i+N}$ are derived from the same image, they are referred to as a positive pair. Then, $\tilde{\bm{x}}_{i}$ is input to $f(\cdot)$, which outputs the vector $\bm{h}_{\tilde{\bm{x}}_{i}}=f(\tilde{\bm{x}}_{i})$. A projection head $g(\cdot)$ consisting of a two-layer multi-layer perceptron (MLP) follows after $f(\cdot)$, and we input $\bm{h}_{\tilde{\bm{x}}_{i}}$ and output vector $\bm{z}_{\tilde{\bm{x}}_{i}}=g(\bm{h}_{\tilde{\bm{x}}_{i}})$. A mini-batch consists of $N$ samples, and data augmentation creates pairs, resulting in a total of $2N$ samples. In contrastive learning, one positive pair $[\bm{z}_{\tilde{\bm{x}}_{i}},\bm{z}_{\tilde{\bm{x}}_{i+N}}]$ and the other $2(N-1)$ negative pairs are consisted, and the positive pairs are attracted and the negative ones are pulled away from each other. Specifically, we perform contrastive learning based on the following InfoNCE loss [59]:
@@
Intra-personal Learning: We acquire feature representations of intra-personal facial changes. The synthetic image $\bm{y}_{i}$ is converted into the image $\tilde{\bm{y}}_{i}$ by data augmentation. Then, $\tilde{\bm{y}}_{i}$ is input to $f(\cdot)$, which outputs the vector $\bm{h}_{\tilde{\bm{y}}_{i}}=f(\tilde{\bm{y}}_{i})$. We next calculate the difference vector $\bm{h}_{\tilde{\bm{y}}_{i}}-\bm{h}_{\tilde{\bm{x}}_{i}}$ between $\bm{h}_{\tilde{\bm{x}}_{i}}$ and $\bm{h}_{\tilde{\bm{y}}_{i}}$ to capture intra-personal facial changes. Furthermore, the difference vector is input to ReLU and a linear layer $l(\cdot)$, which outputs the distance of intra-personal facial changes $d_{i}$. ComFace learns how much the intra-personal face has changed on the basis of the mean squared error (MSE) loss: $\mathcal{L}_{intra}=\frac{1}{N}\sum_{i=1}^{N}(d_{i}-|\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}|)^{2}$, where $|\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}|$ is the ground truth (GT) for the distance of intra-personal facial changes. Since the direction of facial change depends on the attribute, we use $|\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}|$ instead of $\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}$ as the intra-personal GT.
@@ 3.3 Curriculum Learning of Intra-personal Facial Changes
In intra-personal learning, the difficulty level of learning tasks depends on $\alpha$. Specifically, when $\alpha=0.0$ and $5.0$, there are significant differences in the facial changes, so it is easy to identify them (see the left side of Fig. 1). On the other hand, when $\alpha=0.0$ and $2.0$, the facial changes are small, so it is difficult to identify them (see the left side of Fig. 1). ComFace provides effective representation learning by increasing the difficulty level of learning tasks on the basis of curriculum learning. Curriculum learning [5] mimics the human learning behavior of starting with simple tasks and gradually learning more complex concepts. ComFace gradually decreases the distance between intra-personal facial changes $|\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}|$ during training to acquire more effective feature representations. Specifically, suppose that $S$ is the range of the distance of facial changes $|\alpha_{\bm{y}_{i}}-\alpha_{\bm{x}_{i}}|$ when sampling the pairs of $\bm{x}_{i}$ and $\bm{y}_{i}$ used for training, and we gradually decrease $S$ according to the number of epochs as follows:
@@ 4.1 Setup for FRL
Synthetic Face Images: We utilize synthetic face images generated by StyleGAN [29] and StyleGAN3 [28]. The reason for using two different StyleGANs is to employ the attributes for face manipulation provided in each StyleGAN. In StyleGAN, we use the attributes that vary weight provided by Ref. [49] and the attributes that vary age and smile provided by InterFaceGAN [52]. In StyleGAN3, we employ the 40 attributes included in the CelebA dataset [41] provided by Ref. [4]. The 40 attributes contain various face-related factors such as Big Nose, Bags Under Eyes, and Pale Skin. With these 43 attributes, we use synthetic face images that change according to a wide range of face attributes. We generate synthetic face images with 250,000 identities using StyleGANs. For each identity, synthetic images for each attribute with $\alpha=\{-5.0,-4.9,\cdots,0.0,\cdots,4.9,5.0\}$ are generated (see Section 3.1). As a result, the total number of synthetic face images used for FRL is 35 million (M). For details on the composition of synthetic images, please refer to Supplement A.1. See Supplement C.1 for a quality assessment of the synthetic images and an evaluation of the transfer performance with respect to the quality.
Details for FRL: For FRL with synthetic images, we use 90% of all identities as training data and 10% as validation data. As a backbone $f(\cdot)$, we use ResNet50 [23], which is commonly used for general pre-training [13, 7] and representation learning [9, 22, 6]. Since ResNet50 is better than ViT [17] for ComFace (see Supplement C.2 for the evaluation), we employ ResNet50. Synthetic images are resized to 224$\times$224 and the temperature parameter $\tau$ is set to 0.1. Our model is trained from scratch with randomly initialized weights. We run the training for 12 epochs with batch size 1024 on 32 NVIDIA A100 GPUs ($\sim$22 hours training). As parameters for curriculum learning, we set $S_{max}=10,\;t\in\{2,3,4\},\;e_{1}=3,\;e_{2}=6,\;e_{3}=9,\;e_{4}=12$. For each epoch, we randomly sample $\bm{x}_{i}$ and $\bm{y}_{i}$ pairs according to the range $S$ in curriculum learning and construct mini-batches. Adam [32] optimizer is used, and the learning rate is initialized as 4e-4 and halved in 10 epochs. See Supplement B.1 for other setups.
@@ 4.2 Setup for Downstream Tasks
Facial Expression Change Dataset: We use the public dataset DISFA [43, 42], which contains facial videos of 27 subjects (12 women and 15 men) while they watch a 4-minute video intended to elicit a range of facial expressions. For each subject, 4845 video frames were recorded, and the action unit (AU) intensity was annotated for each frame with six levels from 0 (not present) to 5 (maximum intensity) for several AUs. We use AU6 (Cheek Raiser) and AU12 (Lip Corner Puller), which contain frames with a high AU intensity, and we extract frames with an intensity from 1 to 5. We estimate the intensity changes of AU6 and AU12 as facial expression changes. See Supplement C.3 for an evaluation for other major AUs.
@@
See Supplement A.2 for more information on the above datasets.
Details for Downstream Tasks: For each dataset, we perform a four-fold cross-validation that splits subjects between the training and test data. Hence, we evaluate the generalization performance for new subjects not included in the training data. We use 10% of the subjects in the training data as validation data. For both training and testing, the estimation of intra-personal facial changes is performed using two randomly sampled face images within an individual (i.e., approximately the same number of samples in the direction of increasing or decreasing changes). Note that the sampled pairs are identical across methods. For transfer learning, we use two types of evaluation, linear evaluation (i.e., backbone $f(\cdot)$ is frozen and linear layer $l(\cdot)$ is trained from scratch  [^1] See Supplement C.4 for an evaluation in other settings.) and fine-tuning (i.e., all layers in $f(\cdot)$ and $l(\cdot)$ are trained from pre-trained weights  [^2] See Supplement C.5 for an evaluation in other settings.), as in the previous representation learning studies [9, 22, 6]. See Supplement B.2 for other setups.
@@ 4.3 Comparative Methods and Evaluation Metrics
We are curious about the following questions: Does ComFace successfully acquire facial representations that capture intra-personal facial changes? How is the transfer performance of representation learning using synthetic images against real images? To answer these questions, we compare ComFace with four types of methods: scratch (random initialization), general pre-training, visual representation learning, and facial representation learning. We use a supervised method using ImageNet [13] and a face recognition method using VGGFace2 [7] for general pre-training. We also use SimCLR [9], MoCo v2 [22, 10], SwAV [8], and Barlow Twins [65] for visual representation learning. In the above methods, we use ResNet50, which is the same backbone as ours for fair comparison. For facial representation learning, we use Bulat et al. [6], FaRL [70], and PCL [40]. Note that all comparative methods use real images to learn feature representations. In Supplement B.3, we summarize the training datasets, training scales, training sources, and backbones for all comparative and proposed methods.
@@ 4.4 Main Results
Age Change: Table 4 shows the results of estimating age change in fine-tuning. The best models from the four types of comparative methods are listed (see Supplement C.7 for full versions). The table indicates that ComFace has a higher transfer performance for age change than all other methods. We find that our method can be successfully adapted to the task related to long time periods (1 $\sim$) of temporal changes in the face.